A flat silver strip of width 1.5cm and thickness 1.5 mm carries a current of 150 A . A magnetic field of 2T is

applied perpendicular to the flat face of the strip. The emf developed across the width of the strip is measured

to be 17.9 μV. Calculate the free electron density in the silver


1
Expert's answer
2021-08-08T17:14:13-0400

Gives

t=1.5cm

B=2T

i=150A

V=17.9"\mu V"

d=1.5mm

"V_H=\frac{iB}{ned}"

"n=\frac{iB}{V_Hed}"

Put value

"n=\frac{150\times2}{17.9\times10^{-6}\times1.6\times10^{-19}\times1.5\times10^{-3}}"

"n=6.98\times10^{16}" /m3